OpenStudy (anonymous):

y varies jointly as x and the inverse of z. When y = 6 and x = 3, z = 6. Find the constant of variation and determine the value for y when x = 6 and z = 8.

OpenStudy (johnweldon1993):

Back so soon? ;P

OpenStudy (anonymous):

lol yeahhh

OpenStudy (anonymous):

Suppose z varies directly as x and y and inversely as w, and z =8 when x =6, y =5 and w =10. Find z when x =3, y =4 and w = 5. FORMULA: product of "directly" or "jointly" variables "Varies as" variable = k* -------------------------------------------- product of "inversely" variables The "varies as" variable is z, so write: z = k*----- The "directly variables" are x and y, so write them on the top: xy z = k*----- There is just one "inversely variable", w, so write it on the bottom: xy z = k*----- w Now we're ready to substitute the numbers for all the letters except the constant k from this: "z =8 when x =6, y =5 and w =10". Substituting: 6*5 8 = k*----- 10 Solve for k: 30 8 = k*----- 10 8 = k*3 8 --- = k 3 Now go back to this equation: xy z = k*----- w and substitute 3/8 for the constant k. 3 xy z = ---*---- 8 w or 3xy z = ----- 8w Now we substitute in x =3, y =4 and w = 5, to find z: 3(3)(4) z = -------- 8(5) Then finish it up: 36 z = ---- 40 Reduce the fraction: 9 z = ---- 10 The answer is 9/10.

OpenStudy (anonymous):

Thanks ^-^

OpenStudy (anonymous):

np

OpenStudy (anonymous):

wait i dont think thats right tho...

OpenStudy (johnweldon1993):

That's because it isn't ^_^

OpenStudy (johnweldon1993):

y varies jointly as x and the inverse of z. When y = 6 and x = 3, z = 6. Find the constant of variation and determine the value for y when x = 6 and z = 8 \[\large y = \frac{kx}{z}\] "When y = 6 and x = 3, z = 6. Find the constant of variation" \[\large 6 = \frac{3k}{6}\]solve for 'k' here

OpenStudy (anonymous):

thats wat my nots said and i found that on a math site

OpenStudy (johnweldon1993):

@conley1 no that is the correct approach....however they way you have it written with the "9/10 is your answer" made sierra think that was truely HER answer... however @sierraleone17 as you can see...the example @conley1 has given does not have the same numbers as your question nor does it have the same variables

OpenStudy (anonymous):

oh ok